Stuck in the middle of a frozen pond with only your physics book, you decide to put physics in action and throw the 5.00-kg book. If your mass is 62.0 kg and you throw the book at 13.0 m/s, how fast do you then slide across the ice? (Assume the absence of friction.)

Please tell me the explanation of the answer.

1
Expert's answer
2022-03-13T18:46:24-0400

Answer

Using momentum conservation

"mv=m'v'\\v'=\frac{mv}{m'}\\v'=\frac{5*19}{62}=1.53m/s"